in 1995 the life span of males in a certain country was 64.7 years. in 2001 it was 66.7 years. Let E represent the life span in year t represent the number of years since 1995.
E(t)=____t+____
E(11)=

Answer by ankor@dixie-net.com(6691)

in 1995 the life span of males in a certain country was 64.7 years. in 2001 it was 66.7 years. Let E represent the life span in year t represent the number of years since 1995.
:
Find the slope
In 1995; x1=0 and y1=64.7
In 2001; x2=6 and y2=66.7
:
Find the slope (m) using the slope equation: m =
m = = =
therefore:
The Linear function E(t) that fits the data is.
E(t)= t + 64.7
:
E(11)= (11) + 64.7
E(11)= 3.67 + 64.7
E(11)= 68.4 yrs is the life expectancy in 2006